# Soft spots of net negative topological charge directly cause the plasticity of 3D glasses

### v2

@@ Abstract
The deformation mechanism in amorphous solids subjected to external shear remains poorly understood because of the absence of well-defined topological defects mediating the plastic deformation. The notion of soft spots has emerged as a useful tool to characterize the onset of irreversible rearrangements and plastic flow, but these entities are not well-defined in terms of geometry and topology. In this study, we unveil the phenomenology of recently discovered, well-defined topological defects governing the microscopic mechanical and yielding behavior of a model 3D glass under shear deformation. We identify the existence of vortex-like and anti-vortex-like topological defects within the 3D non-affine displacement field. The number density of these defects exhibits a significant (inverse) correlation with the plastic events, with defect proliferation-annihilation cycles matching the alternation of elastic-like segments and catastrophic plastic drops, respectively. Furthermore, we observe collective annihilation of these point-like defects via plastic events, with large local topological charge fluctuations in the vicinity of regions that feature strong non-affine displacements. We unveil that plastic yielding is driven by very few, but very large, clusters of net negative topological charge, the massive annihilation of which triggers the onset of plastic flow. These findings suggest a geometric and topological characterization of soft spots and pave the way for the mechanistic understanding of topological defects as mediators of plastic deformation in glassy materials.
@@
Introduction – While plastic deformations in crystals are by now well understood in terms of dislocation dynamics, a concept originally proposed by Taylor Taylor (1934), identifying the physical mechanism of plasticity in amorphous solids, such as glasses, is more challenging Wang et al. (2015). Despite several attempts to generalize the concept of topological defects and dislocation-like structures in amorphous solids Gilman (2003); Chaudhari et al. (1979); Steinhardt and Chaudhari (1981); Egami et al. (1980), not much progress in this direction has been achieved so far, and this approach is still considered by many as hopeless, due to the unavoidable disordered reference structure Alexander (1998). Other existing viewpoints on this matter, such as shear transformation zone (STZ) theory Falk and Langer (1998), 2D melting-type scenarios Fraggedakis et al. (2023) or elastoplastic models Nicolas et al. (2018), assume the existence of soft spots linked to some unspecified defects, but they are purely phenomenological and agnostic about the precise definition of the latter. In particular, soft spots, as discussed in the literature Manning and Liu (2011), lack a quantitative geometric characterization or measure to identify them, in contrast to topological defects in crystalline matter and liquid crystals Fumeron and Berche (2023). At the same time, novel ideas to identify defects in amorphous solids based on quasi-localized modes Lerner and Bouchbinder (2021), geometric charges Lemaître et al. (2021) or string-like defects Lund (2015); Zhou et al. (2023) are not mature enough; consequently, structural indicators Richard et al. (2020) remain still the best available option to predict plasticity in glasses.
@@
Almost at the same time with the current paper, a preprint Desmarchelier et al. (2024) has been posted, where the concept of topological defects has been used to unambiguously identify shear transformation zones in 2D amorphous solids, providing a direct connection between individual topological defects with negative charge and local structural rearrangements responsible for plasticity.
In this study, we establish a protocol for identifying quantized (±$\pm 1$) topological defects in the displacement field of model 3D glasses undergoing shear deformation. Through defect analysis of diverse 2D slices within the shear plane, we demonstrate a correlation between the population of topological defects and plastic drops in stress-strain. Further dissection of these 2D slices reveals microscopic details on local topological charge and associated non-affinity. Notably, fluctuations in topological charge are observed in proximity to regions exhibiting strong non-affine displacements, indicating the emergence of soft spots. Systematic analysis of various 2D slices enables the construction of a three-dimensional representation of topological charges originating from point defects, highlighting the drastic annihilation of such defects during yielding – a “topological catastrophe”.
Results – Here we prepare a polymer glass sample at $T=0$ implemented using the Kremer-Grest model Kremer and Grest (1986). The system is then subjected to an athermal quasi-static (AQS) shearing protocol involving affine transformation by small strain increments of $\delta\gamma_{xz}=0.001$ in the $xz$ plane following energy minimization. More details on the interaction potential and the deformation protocol of the simulations are provided in the method section.
@@
The non-affine field is defined as ${\bf{u}}={\bf{u_{t}}}-{\bf{u_{A}}}$ Zaccone (2023), and can be expressed as ${\bf{u}}=\bf{u_{\parallel}}+\bf{u_{\perp}}$, where $\bf{u_{\parallel}}$ is the vector in the $xz$ (shear) plane, and $\bf{u_{\perp}}$ is the vector along the $y$-direction. The average magnitude of non-affine displacement in three dimensions is computed as $u_{\rm{mag}}^{3D}=\frac{1}{N}\sum_{i=1}^{N}|\bf{u}_{i}|$. As the shear is applied along the $x$-direction to deform the $xz$ (shear) plane, we compute the average non-affine displacement in the $xz$ plane as $u_{\rm{mag}}=\langle\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}|\bf{u_{{\parallel},i}}|\rangle$, where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over all slices along the $y$ direction, and $N_{s}$ represents the number of particles in each two-dimensional (2D) slice. For our analysis, we divided the 3D box into numerous two-dimensional slices of width $2\sigma$ ($\sigma$ is the monomer diameter) along the $y$ direction.
@@
In Fig.1, we present the stress ($\sigma$) versus strain ($\gamma$) curve. Additionally, we plot the average magnitudes of the non-affine displacement, denoted as $u_{\rm{mag}}^{3D}$ and $u_{\rm{mag}}$, represented by red and blue lines, respectively. The stress drop is notably identifiable by the peaks in the non-affine displacement values. These peaks in $u_{\rm{mag}}^{3D}$ and $u_{\rm{mag}}$ are evidently commensurate with the occurrence of plastic events. We have highlighted several of these plastic events using green dashed lines in Fig.1. Remarkably, we observe that $u_{\rm{mag}}$, which disregards the $y$-component of the field, can accurately identify the locations of plastic events. This observation implies that the characteristics of displacement field remain consistent when we exclude the $y$-component (orthogonal to the shear plane).
@@
In 2D, a point defect carrying a winding number of $q=+1$ is recognized as a vortex, while a defect with $q=-1$ is identified as an anti-vortex. We systematically discerned these point defects across various 2D slices within the $xz$ plane. As depicted in Fig.1(b), we present the interpolated non-affine displacement field with a color bar indicating the scaled magnitude, within a 2D slice at $y=0$ for $\gamma=0.098$. The existence of topological defects with winding numbers $q=+1$ and $q=-1$ is marked by magenta and black-filled circles, respectively. Notably, two rectangular regions are selectively magnified to elucidate the chirality of vortices ($q=+1$). At the frustrated interface between two vortices of identical phase chirality (direction of swirls), an anti-vortex ($q=-1$) is observed. Conversely, vortices with opposing phase chiralities exhibit a stable interface, as portrayed in the enlarged sections of Fig. 1(b). This intriguing pattern bears resemblance to topological defects discerned in the eigenvectors field of a recently investigated 2D glass Wu et al. (2023). As shown in Desmarchelier et al. (2024), the displacement and stress distribution of $-1$ defects strongly resemble those of an Eshelby inclusion with quadrupolar structure, i.e. a shear transformation zone.
@@
In Fig.2, we present the locations of point-defects within the $xz$ plane non-affine displacement field at $y=0$ for four consecutive $\gamma$ values that straddle the yielding point. These snapshots reveal a significant transformation in the structure of defects. We compute the total number of these point defects in the non-affine displacement field across various $xz$ planes by varying $y$ and calculate the average number of defects $N_{d}$ per slice. In order to establish a connection between plastic events and the emergence of point-like topological defects, we compute $N_{d}$ with increasing strain $\gamma$.
In Fig.3(a), we illustrate the relative changes in the average magnitude of the displacement field ($u_{\rm{mag}}$) and the average number of defects ($N_{d}$) per slice within a range of $\gamma$. It is notable that the peaks in $u_{\rm{mag}}$ correspond to dips in the total number of point defects. This observation strongly suggests an anti-correlation between $u_{\rm{mag}}$ and $N_{d}$, implying that accumulation of topological defects is the trigger of plastic events via subsequent defects annihilation.
To explore this dependence further, we calculate the Pearson correlation coefficient (PCC) Pearson and Galton (1895) as:
${\rm{PCC}}=\frac{\sum_{\{\gamma\}}({u}_{\rm{mag}}-\bar{u}_{\rm{mag}})(N_{d}-\bar{N}_{d})}{\sqrt{\sum_{\{\gamma\}}({u}_{\rm{mag}}-\bar{u}_{\rm{mag}})^{2}}\sqrt{\sum_{\{\gamma\}}(N_{d}-\bar{N}_{d})^{2}}},$ (2)
@@
where $\bar{u}_{\rm{mag}}$ and $\bar{N}_{d}$ represent the average values of $u_{\rm{mag}}$ and $N_{d}$ over the range of strain $\gamma\in[0,0.2]$, respectively. The value of ${\rm{PCC}}$ falls within the range of [-1, 1], where a value close to 1 indicates a strong positive correlation, while a value near -1 suggests a strong negative correlation. For the considered range of $\gamma$, we obtain the value of ${\rm{PCC}}\approx-0.5$, indicating a marked level of anti-correlation. Interestingly, when we consider only the data points associated with plastic events, identified from the peaks of $u_{\rm{mag}}$, the value of ${\rm{PCC}}$ approaches $-1$. This suggests that there is a strong (inverse) relationship between the topological defects and plastic deformation events in the material.
@@
Figure 3: (a) The magnitude of non-affine displacement field ($u_{\rm{mag}}$) and the average defect number $N_{d}$ are shown in certain range of $\gamma$ values. Both quantities are presented after subtracting their respective average values, and the $N_{d}$ values have been scaled by 0.01 for better visualization. (b) A 2D slice ($\gamma=0.099$ and $y=0$) dissected into many unit area square cells is shown with the local topological number density $\rho_{d}$ in each cell by color bar. The cells having larger non-affine magnitude have been presented by red crosses. (c) The same slice, as in (b), shown by color (scaled by 0.2) which represents the local magnitude of non-affine displacement. The cells marked by crosses colored in magenta and black are populated by topological charges with positive and negative values, respectively. (d) The average topological charge density $\bar{C}_{avg}$ per slice is shown with the variation of $\gamma$.
@@
While we have established a robust anti-correlation between plastic events and the variation in the total number of topological point defects, our primary objective is to understand the microscopic details of these plastic events and investigate the specific role played by the topological defects. Upon examining the snapshots in Figs.1(b) and 2, we observe that topological defects tend to be positioned in close proximity to regions characterized by high non-affine displacement magnitudes. In order to investigate whether the local non-affine displacement field correlates with the presence of topological defects, we systematically partition each 2D slice into numerous unit area square cells, as illustrated in Figs.3(b) and (c). Within each such cell, we compute the number density of defects $n_{d}$, average non-affine displacement $u_{\rm{mag}}$ and concurrently measure the local topological charge in each cell, denoted as $\bar{C}_{k}=\sum_{j\in\mathbb{S}_{k}}q_{j}$, where $q_{j}$ is the topological charge of defects present inside the $k$-th cell, $\mathbb{S}_{k}$. In fact, $\bar{C}_{k}$ essentially measures the topological charge imbalance inside each cell. The average topological charge imbalance is computed as $\bar{C}_{avg}=(1/n_{\rm{cell}})\sum_{k=1}^{n_{\rm{cell}}}\bar{C}_{k}$, where $n_{\rm{cell}}$ is the total number of square cells. Let us emphasize that this analysis probes a mesoscopic coarse-grained length scale corresponding to the size of the (square) sub-cell of length $\sigma$ and does not reveal any local microscopic information.
In Fig.3(b) and (c), we visually illustrate the partitioning of a 2D slice into multiple square cells, utilizing a color bar to represent the number density of topological defects ($n_{d}$) and the magnitude of the non-affine displacement field ($u_{\rm{mag}}$ within each cell), respectively. The red crosses in Fig. 3(b) pinpoint cells where $u_{\rm{mag}}$ is notably higher compared to other cells, while colours of the crosses within several cells in Fig.3(c) demarcate the signs of the average topological charge, denoted as $\bar{C}_{avg}$. Magenta highlights cells with $\bar{C}_{k}\geq 1$, and black designates cells with $\bar{C}_{k}\leq-1$. In Fig.3(d), we present the average local topological charge imbalance $\bar{C}_{avg}$, computed across all cells and various 2D slices, while varying $\gamma$. Remarkably, we observe a prevailing bias in local charge imbalance towards net negative charge values, which we attribute as an underlying tendency towards plastic deformation.
@@
Our analysis has primarily focused on various 2D cross-sections within the $xz$ plane along the $y$ axis. In order to extend our observations into three dimensions, we developed a generalized protocol to systematically scan the system from arbitrary directions. Positions of topological defects were identified across numerous two-dimensional slices, each oriented along a direction vector $\hat{n}$. By considering various directions $\hat{n}$ uniformly distributed on a unit sphere, we located defects within the simulation box. Similar to the dissection of 2D slices illustrated in Fig. 3(b) and (c), we partitioned the entire simulation box into cubical cells of unit volume. Utilizing the locations of topological defect points identified from 2D dissections along all possible directions, we computed the local topological charge imbalance $\bar{C}^{c}_{k}$ (scaled within the range [0,1]) within these cubical cells.

### v3

# Clustering of negative topological charge precedes plastic failure in 3D glasses

@@ Abstract
The deformation mechanism in amorphous solids subjected to external shear remains poorly understood because of the absence of well-defined topological defects mediating the plastic deformation. The notion of soft spots has emerged as a useful tool to characterize the onset of irreversible rearrangements and plastic flow, but these entities are not clearly defined in terms of geometry and topology. In this study, we unveil the phenomenology of recently discovered, precisely defined topological defects governing the microscopic mechanical and yielding behavior of a model 3D glass under shear deformation. We identify the existence of vortex-like and anti-vortex-like topological defects within the 3D non-affine displacement field. The number density of these defects exhibits a significant anti-correlation with the plastic events, with defect proliferation-annihilation cycles matching the alternation of elastic-like segments and catastrophic plastic drops, respectively. Furthermore, we observe collective annihilation of these point-like defects via plastic events, with large local topological charge fluctuations in the vicinity of regions that feature strong non-affine displacements. We reveal that plastic yielding is driven by several large sized clusters of net negative topological charge, the massive annihilation of which triggers the onset of plastic flow. These findings suggest a geometric and topological characterization of soft spots and pave the way for the mechanistic understanding of topological defects as mediators of plastic deformation in glassy materials.
@@
While plastic deformations in crystals are by now well understood in terms of dislocation dynamics, a concept originally proposed by Taylor Taylor (1934), identifying the physical mechanism of plasticity in amorphous solids, such as glasses, is more challenging Wang et al. (2015). Despite several attempts to generalize the concept of topological defects and dislocation-like structures in amorphous solids Gilman (2003); Chaudhari et al. (1979); Steinhardt and Chaudhari (1981); Egami et al. (1980), not much progress in this direction has been achieved so far, and this approach is still considered by many as hopeless. The deformed structure is indistinguishable from the disordered reference structure Alexander (1998). Other existing viewpoints on this matter, such as shear transformation zone (STZ) theory Falk and Langer (1998), 2D melting-type scenarios Fraggedakis et al. (2023) or elastoplastic models Nicolas et al. (2018), assume the existence of soft spots linked to some unspecified defects, but they are purely phenomenological and agnostic about the precise definition of the defect. In particular, soft spots, as discussed in the literature Manning and Liu (2011), lack a quantitative geometric characterization or measure to identify them, in contrast to topological defects in crystalline matter and liquid crystals Fumeron and Berche (2023). At the same time, novel ideas to identify defects in amorphous solids based on quasi-localized modes Lerner and Bouchbinder (2021), geometric charges Lemaître et al. (2021) or string-like defects Lund (2015); Zhou et al. (2023) are not mature enough; consequently, structural indicators Richard et al. (2020) still remain the best available option to predict plasticity in glasses.
@@
Almost at the same time with the current paper, a preprint Desmarchelier et al. (2024) has been posted, where the concept of topological defects has been used to unambiguously identify shear transformation zones in 2D amorphous solids, providing a direct connection between individual topological defects with negative charge and local structural rearrangements responsible for plasticity. Topological defects can be modelled in a continuum fashion using geometry, and in particular relating dislocations to torsion and disclinations to curvature Kleinert (1989); Kleman and Friedel (2008). Likewise, point defects can be described using a geometric approach Kupferman et al. (2018); Yavari and Goriely (2012). The relationship between these geometric charges and the vortex-like defects considered in this article remains unclear. Our work investigates the connection between topological defects in the non-affine displacement field, as examined in recent studies Wu et al. (2023); Desmarchelier et al. (2024), and the mechanical properties of a three dimensional polymer glass system under shear deformation.
In this study, we establish a protocol for identifying quantized (±$\pm 1$) topological defects in the displacement field of model 3D glasses undergoing shear deformation. Through defect analysis of quasi-2D slices within the shear plane, we demonstrate a correlation between the population of topological defects and plastic drops in the stress-strain curve. Further dissection of these 2D slices reveals microscopic details on local topological charge and associated non-affinity. Notably, fluctuations in topological charge are observed in proximity to regions exhibiting strong non-affine displacements, indicating the emergence of soft spots. Systematic analysis of various 2D slices enables the construction of a three-dimensional representation of topological charges originating from defects, highlighting the drastic annihilation of such defects during yielding – a “topological avalanche” Altshuler and Johansen (2004).
Results
Here we prepare a polymer glass sample at $T=0$ implemented using the Kremer-Grest model Kremer and Grest (1986). The system is then subjected to an athermal quasi-static (AQS) shearing protocol involving affine transformation by small strain increments of $\delta\gamma_{xz}=0.001$ in the $xz$ plane following energy minimization. More details on the interaction potential and the deformation protocol of the simulations are provided in the method section.
@@ Results
The non-affine field is defined as ${\bf{u}}={\bf{u_{t}}}-{\bf{u_{A}}}$ Zaccone (2023), and can be expressed as ${\bf{u}}=\bf{u_{\parallel}}+\bf{u_{\perp}}$, where $\bf{u_{\parallel}}$ is the vector in the $xz$ (shear) plane, and $\bf{u_{\perp}}$ is the vector along the $y$-direction. The average magnitude of non-affine displacement in three dimensions is computed as $u_{\rm{mag}}^{3D}=\frac{1}{N}\sum_{i=1}^{N}|\bf{u}_{i}|$. This average is performed over the total number of monomers/particles ($N=10000$) in the system. As the shear is applied along the $x$-direction to deform the $xz$ (shear) plane, we compute the average non-affine displacement in the $xz$ plane as $u_{\rm{mag}}=\langle\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}|\bf{u_{{\parallel},i}}|\rangle$, where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over all slices along the $y$ direction, and $N_{s}$ represents the number of particles in each two-dimensional (2D) slice. For our analysis, we divided the 3D box into numerous two-dimensional slices of thickness $2\sigma$ ($\sigma$ is the monomer diameter) along the $y$ direction.
@@
In Fig.1, we present the stress ($\sigma$) versus strain ($\gamma$) curve. Additionally, we plot the average magnitudes of the non-affine displacement, denoted as $u_{\rm{mag}}^{3D}$ and $u_{\rm{mag}}$, represented by red and blue lines, respectively. The stress drop is notably identifiable by the peaks in the non-affine displacement values. These peaks in $u_{\rm{mag}}^{3D}$ and $u_{\rm{mag}}$ are evidently commensurate with the occurrence of plastic events (PEs). We have highlighted several of these PEs using green dashed lines in Fig.1. Remarkably, we observe that $u_{\rm{mag}}$, which disregards the $y$-component of the field, can accurately identify the locations of PEs. This observation implies that the characteristics of displacement field remain consistent when we exclude the $y$-component (orthogonal to the shear plane). However, the same behavior is observed even when we exclude the $x$ or $z$ component of the non-affine field, as proven explicitly in the Supplementary Information. This suggests that the peaks in $u_{\rm{mag}}$ occur due to the high values of the non-affine field magnitude for a small fraction of all particles. All three components of the non-affine displacement for these fraction of particles have approximately of the same order of magnitude. This is why discarding any one component does not alter the overall characteristics of the $u_{\rm{mag}}$ versus $\gamma$ curve.
Moreover, we have confirmed the convergence of our numerical analysis by considering also smaller strain steps, $\delta\gamma=10^{-4}$, and obtained analogous results. As a direct proof, we compare the magnitude of non-affine displacement for different strain steps in the Supplementary Information.
@@
In 2D, a point defect carrying a winding number of $q=+1$ is recognized as a vortex, while a defect with $q=-1$ is identified as an anti-vortex. We systematically discerned these point defects across various 2D slices within the $xz$ plane. As depicted in Fig.1(b), we present the interpolated non-affine displacement field with a color bar indicating the scaled magnitude, within a 2D slice at $y=0$ for $\gamma=0.098$. The existence of topological defects with winding numbers $q=+1$ and $q=-1$ is marked by magenta and black-filled circles, respectively. Notably, two rectangular regions are selectively magnified to elucidate the chirality of vortices ($q=+1$). At the frustrated interface between two vortices of identical phase chirality (direction of swirls), an anti-vortex ($q=-1$) is observed. Conversely, vortices with opposing phase chiralities exhibit a stable interface, as portrayed in the enlarged sections of Fig. 1(b). This intriguing pattern bears resemblance to topological defects discerned in the eigenvectors field of a recently investigated 2D glass Wu et al. (2023). As shown in Desmarchelier et al. (2024), the displacement and stress distribution of $-1$ defects strongly resemble those of an Eshelby inclusion with quadrupolar structure, i.e. a shear transformation zone. The non-affine displacement field incorporates contributions from the entire spectrum of eigenmodes of the Hessian matrix. In the work by Tanguy Tanguy et al. (2002), it is reported that the structure of the eigenvectors from the Hessian matrix and that from the nonaffine displacement field strongly correlate in space for low eigenfrequencies. Since plastic events are typically dominated by low-frequency modes, it can be expected that analyses based on these low-frequency modes and the displacement field yield consistent results. However, this aspect is beyond the scope of our current study.
@@
In Fig.2, we present the locations of point-defects within the $xz$ plane non-affine displacement field at $y=0$ for four consecutive $\gamma$ values that straddle the yielding point. These snapshots reveal a significant transformation in the structure of defects. We compute the total number of these point defects in the non-affine displacement field across various $xz$ planes by varying $y$ and calculate the average number of defects $N_{d}$ per slice. In order to establish a connection between plastic events and the emergence of point-like topological defects, we compute $N_{d}$ with increasing strain $\gamma$. In our analysis we interpolate the displacement field, ignoring the periodic boundary conditions, and additionally, we consider two dimensional slices of a 3D system. Consequently, the number of vortices and anti-vortices is not completely equal in any given interpolated 2D displacement field. However, when averaging the defect numbers over several 2D slices, the number of defects with opposite charge is approximately equal; leading to near-zero total topological charge, or approximate conservation of topological charge.
In Fig.3(a), we depict the average magnitude of the displacement field ($u_{\rm{mag}}$) and the average number of defects ($N_{d}$) per slice across the entire range of $\gamma$. To establish a correlation, we calculate the dimensionless quantities $\tilde{u}_{\rm{mag}}$ and $\tilde{N}_{d}$, defined as $\tilde{u}_{\rm{mag}}=({u}_{\rm{mag}}-\bar{u}_{\rm{mag}})/{{\sigma}_{u_{\rm{mag}}}}$ and $\tilde{N}_{d}=({N}_{d}-\bar{N}_{d})/{{\sigma}_{N_{d}}}$. Here $\bar{u}_{\rm{mag}}$ and $\bar{N}_{d}$ denote the average values of $u_{\rm{mag}}$ and $N_{d}$ over the range of strain $\gamma\in[0,0.2]$, respectively. The quantities ${{\sigma}^{2}_{u_{\rm{mag}}}}={\sum_{\{\gamma\}}({u}_{\rm{mag}}-\bar{u}_{\rm{mag}})^{2}}$ and ${{\sigma}^{2}_{N_{d}}}={\sum_{\{\gamma\}}({N}_{d}-\bar{N}_{d})^{2}}$ are proportional to the variances of $u_{\rm{mag}}$ and $N_{d}$, respectively. In Fig.3(b), we illustrate the variation of $\tilde{u}_{\rm{mag}}$ and $\tilde{N}_{d}$ within the range of $\gamma$. Notably, the peaks in $u_{\rm{mag}}$ align with dips in the total number of point defects. This observation strongly suggests an anti-correlation between $u_{\rm{mag}}$ and $N_{d}$, implying that the accumulation of topological defects triggers plastic events through subsequent defect annihilation. This anti-correlation might be explained by the fact that topological defects occur where the displacement field is close to zero and fluctuating, and thus singular or not defined. It is thus plausible that these defects anti-correlate with plasticity, as zeros in the displacement field are less common at strain steps pertaining to large plastic events, concomitant with large and coordinated displacements over extensive regions.
As apparent from Fig.3(a), the topological defects are of transient nature. In the Supplementary Information, we have analyzed this aspect further by extending the analysis to smaller strain step, $\delta\gamma=10^{-4}$. We observe drastic fluctuations in $N_{d}$ in this case as well, indicating that the transient nature persists even for smaller $\delta\gamma$. In order to understand the origin of this transient nature, it would be useful to investigate in detail the dynamics of these defects, their lifetime and their interactions.
@@
To explore this dependence further, we calculate the Pearson correlation coefficient (${\rm{PCC}}$) Pearson and Galton (1895), given by ${\rm{PCC}}=\sum_{\{\gamma\}}\tilde{u}_{\rm{mag}}\tilde{N}_{d}$. The value of ${\rm{PCC}}$ falls within the range of [-1, 1], where a value close to 1 indicates a strong positive correlation, while a value near -1 suggests a strong negative correlation. Averaging over the entire range of $\gamma$, we obtain the value of ${\rm{PCC}}\approx-0.5$, indicating a marked level of anti-correlation. Interestingly, when we consider only the data points associated with plastic events, identified from the peaks of $u_{\rm{mag}}$, the value of ${\rm{PCC}}$ approaches $-1$. This suggests that there is a strong (inverse) relationship between the topological defects and plastic deformation events in the material.
Figure 3: (a) The magnitude of non-affine displacement field $u_{\rm{mag}}$ (left) and the average defect number $N_{d}$ (right) are shown for the entire range of $\gamma$ values. (b) The quantities, $\tilde{u}_{\rm{mag}}$ and $\tilde{N}_{d}$ are presented for a limited range of $\gamma$ for better visualization.
@@
Figure 4: (a) A 2D slice ($\gamma=0.099$ and $y=0$) dissected into many unit area square cells is shown with the local topological number density $\rho_{d}$ in each cell by color bar. The cells having larger non-affine magnitude have been presented by red crosses. (b) The same slice, as in (a), shown by color (scaled by 0.2) which represents the local magnitude of non-affine displacement. The cells marked by crosses colored in magenta and black are populated by topological charges with positive and negative values, respectively. (c) The average topological charge density $\bar{C}_{\mathrm{avg}}$ per cell is shown as a function of $\gamma$ for two different coarse-grained sizes of the sub-square box $a_{s}\times a_{s}$.
@@
While we have established a robust anti-correlation between plastic events and the variation in the total number of topological point defects, our primary objective is to understand the microscopic details of these plastic events and investigate the specific role played by the topological defects. Upon examining the snapshots in Figs.1(b) and 2, we observe that topological defects tend to be positioned in close proximity to regions characterized by high non-affine displacement magnitudes. In order to investigate whether the local non-affine displacement field correlates with the presence of topological defects, we systematically partition each 2D slice into numerous unit area square cells, as illustrated in Figs.4(a) and (b). Within each such cell, we compute the number density of defects $n_{d}$, average non-affine displacement $u_{\rm{mag}}$ and concurrently measure the local topological charge in each cell, denoted as $\bar{C}_{k}=\sum_{j\in\mathbb{S}_{k}}q_{j}$, where $q_{j}$ is the topological charge of defects present inside the $k$-th cell, $\mathbb{S}_{k}$. In fact, $\bar{C}_{k}$ essentially measures the topological charge imbalance inside each cell. The average topological charge imbalance is computed as $\bar{C}_{\mathrm{avg}}=(1/n_{\rm{cell}})\sum_{k=1}^{n_{\rm{cell}}}\bar{C}_{k}$, where $n_{\rm{cell}}$ is the total number of square cells. Let us emphasize that this analysis probes a mesoscopic coarse-grained length scale corresponding to the size of the (square) sub-cell of length $\sigma$ and does not reveal any local microscopic information.
In Fig.4(a) and (b), we visually illustrate the partitioning of a 2D slice into multiple square cells, utilizing a color bar to represent the number density of topological defects ($n_{d}$) and the magnitude of the non-affine displacement field ($u_{\rm{mag}}$ within each cell), respectively. The red crosses in Fig. 4(a) pinpoint cells where $u_{\rm{mag}}$ is notably higher compared to other cells, while colours of the crosses within several cells in Fig.4(b) demarcate the signs of the average topological charge, denoted as $\bar{C}_{\rm{avg}}$. Magenta highlights cells with $\bar{C}_{k}\geq 1$, and black designates cells with $\bar{C}_{k}\leq-1$. In Fig.4(c), we present the average local topological charge imbalance $\bar{C}_{\rm{avg}}$, computed across all cells and various 2D slices, while varying $\gamma$. Here, we depict the variation of $\bar{C}_{{\rm{avg}}}$ for two different square cell areas with sides $a_{s}=1$ and $2$. The behavior of $\bar{C}_{{\rm{avg}}}$ versus $\gamma$ remains consistent across varying coarse-grained cell sizes, except for a constant scale factor reflecting the different cell areas. Remarkably, we observe a prevailing bias in local charge imbalance towards net negative charge values, which we attribute as an underlying tendency towards plastic deformation.
@@
Our analysis has primarily focused on various 2D cross-sections within the $xz$ plane along the $y$ axis. In order to extend our observations into three dimensions, we developed a generalized protocol to systematically scan the system from arbitrary directions. Positions of topological defects were identified across numerous two-dimensional slices, each oriented along a direction vector $\hat{n}$. By considering various directions $\hat{n}$ uniformly distributed on a unit sphere, we located defects within the simulation box. Similar to the dissection of 2D slices illustrated in Fig. 4(a) and (b), we partitioned the entire simulation box into cubical cells of unit volume. Utilizing the locations of topological defect points identified from 2D dissections along all possible directions, we computed the local topological charge imbalance $\bar{C}^{c}_{k}$ (scaled within the range [-1,1]) within these cubical cells.
In Fig.7, 3D simulation boxes are depicted, wherein coarse-grained cells with $\bar{C}^{c}_{k}<-0.6$ are highlighted in light blue. Additionally, we identify connected cubical cells with large negative charges, referred to as clusters. In Table 1, we present the number of clusters $N_{cls}$ with a size of $n_{c}\geq 8$ just before a PE, at the PE, and after the PE for various values of $\gamma$. In all cases, we observe a significant decrease in $N_{\text{cls}}$ at PEs, followed by an increase after each PE. Furthermore, Fig.7 (a)-(b) show all $11$ clusters with $n_{c}\geq 8$ just before yielding and all $6$ clusters at the PE. A similar scenario is depicted in Fig.7 (c)-(d) for another PE at $\gamma=0.154$. Notably, we observe the disappearance of most of these large connected clusters at PEs and their subsequent development just after the PEs. This suggests that the clustering of negative topological charges is a precursor of a PE and that PEs can be identified with the sudden fragmentation of these large clusters of negative topological charge.
Table 1: The number of clusters $N_{cls}$ having size $n_{c}\geq 8$ for several plastic events.
Strain $\gamma$ at PE $N_{\text{cls}}$ (before) $N_{\text{cls}}$ (at PE) $N_{\text{cls}}$ (after)
$0.099$ $11$ 6 13
$0.106$ $15$ 4 8
$0.120$ $7$ 4 8
$0.126$ $8$ 4 9
$0.154$ $11$ 3 9